For a science experiment you need to electroplate zinc onto both sides of a very thin, 1.00 cm × 1.00 cm copper sheet. You want the zinc thickness to be 140 nm , and you know that the charge carriers in the ionic solution are divalent (charge 2e) zinc ions. The density of zinc is 7140 kg/m3 If the electroplating apparatus operates at 1.40 mA , how long will it take the zinc to reach the desired thickness?
Answer #1

FIrst we have to find the mass of zinc to be formed

thickness = 140 nm

140 nm \times \frac{10^{-9}m}{1 nm}

=1.40x10-7 m

Surface area of the electrode = 1 x 1 = 1 cm2 =

1cm^{2} \times \frac{(10^{-2}m)^{2}}{(1 cm)^{2}}

=10-4 m2

Volume of zinc = area x thickness = 10-4 m2x1.40x10-7 m = 1.40x 10-11 m3

From density we can calculate mass

mass = density x volume = 7140 kg/m3 x 1.40x 10-11 m3 =1.0x10-6 kg = 1.0x10-3 g

For 1 mole Zn requires 2F of current

moles of Zn =

1.0\times 10^{-3} g Zn\times \frac{1 mol Zn}{65.48 g}

=1.53x10-6 mol Zn

Current required to deposit 1.53x10-6 mol Zn =

1.53\times 10^{-6} mol Zn\times \frac{2 \times 96485 C}{1 mol Zn}

Q = 0.295 C

Formula is Q= it

i = 1.40 mA

time = Q/i = 0.295/1.40x10-3 s

time =210 s